# Exploring Radio Emissions from Confirmed Exoplanets Using SKA

### v1

@@ 2.1 Estimation of Exoplanets Radio Emission
Observations of magnetized planets within our Solar System have revealed a consistent relationship between the power of auroral emissions and the energy flux of the incoming stellar wind (Zarka et al., 2001; Zarka, 2004, 2007; Zarka et al., 2018). This phenomenon extends to various planetary bodies, including those with satellite-Jupiter radio emissions (Noyola et al., 2014, 2016). Specifically, studies have shown that approximately 1% of the energy carried by electrons in these systems manifests as radio waves (Zarka et al., 2018). This correlation between auroral emissions and stellar wind energy flux, along with the strength of the planetary magnetic field, is encapsulated by the RBL model (Zarka et al., 2001; Zarka, 2004, 2007). Mathematically, this model relates the emitted frequency to the magnetic field strength, as expressed by
@@ 3 Results
Figure 1 displays the distributions of planetary parameters, with exoplanets exhibiting detected radio fluxes depicted in distinct colors. It’s noticeable that a significant portion of exoplanets emitting intense radio fluxes do so at frequencies below the sensitivity range of the SKA telescope (below 50 MHz). One method to expand the detection of exoplanets with observable radio signals using current technology involves leveraging microlensing events. These events act as natural amplifiers, their magnification unaffected by the radiation’s wavelength. A notable characteristic of gravitational lensing is the preservation of radiation polarization during the event. Unlike unpolarized stellar plasma radiation, radio emissions from exoplanets exhibit circular polarization (Zarka et al., 2014). Consequently, microlensing events offer a potential means to enhance the detection of exoplanets with observable radio signals (Bagheri et al., 2024).

### v2

@@ 2.1 Estimation of Exoplanets Radio Emission
In this section we follow the methodology outlined in (Bagheri et al., 2024) to estimate the rdaio emission of exoplanets. Observations of magnetized planets within our Solar System have unveiled a consistent link between the intensity of auroral emissions and the energy flux of the incoming stellar wind (Zarka et al., 2001; Zarka, 2004, 2007; Zarka et al., 2018). This phenomenon extends across various planetary bodies, including those exhibiting satellite-Jupiter radio emissions (Noyola et al., 2014, 2016). The correlation between auroral emissions, stellar wind energy flux, and the strength of the planetary magnetic field is encapsulated within the RBL model (citepzarka2001magnetically, zarka2004fast, zarka2007plasma. Mathematically, this model establishes a relationship between the emitted frequency and the magnetic field strength, as outlined by
@@ 3 Results
Figure 1 displays the distributions of planetary parameters, with exoplanets exhibiting detected radio fluxes depicted in distinct colors. It’s noticeable that a significant portion of exoplanets emitting intense radio fluxes do so at frequencies below the sensitivity range of the SKA telescope (below 50 MHz). One method to expand the detection of exoplanets with observable radio signals using current technology involves leveraging microlensing events. These events act as natural amplifiers, their magnification unaffected by the radiation’s wavelength. A notable characteristic of gravitational lensing is the preservation of radiation polarization during the event. Unlike unpolarized stellar plasma radiation, radio emissions from exoplanets exhibit circular polarization (Zarka et al., 2014). Consequently, microlensing events offer a potential means to enhance the detection of exoplanets with observable radio signals (Bagheri et al., 2024, 2019).